# Privacy Guarantees in Posterior Sampling under Contamination

### v2

@@ B.3.1 Proof for Lemma 10
Proof for Lemma 10.
It is enough to check that a $u$-bracket for $\mathcal{K}_{p^{\prime}}(A)$ is also a $u\sqrt{\frac{1-p}{1-p^{\prime}}}$-bracketing for $\mathcal{K}_{p}(A)$, where $1>p\geq p^{\prime}$. Suppose that $\{[L^{\prime}_{i},U^{\prime}_{i}],i=1,\dots,N\}$ is a $u$-bracketing for $\mathcal{K}_{p^{\prime}}(A)$. Pick an arbitrary $\theta\in A$, there exists $i\in\{1,\dots,N\}$ such that the $u$-bracket $[L^{\prime}_{i},U^{\prime}_{i}]$ forms an envelope for $k_{p^{\prime}}(\bm{x};\theta)$, i.e.,
@@
$=$ $\,\frac{1-p}{1-p^{\prime}}u^{2}$
where the second inequality is by noting that for any $a\geq b\geq 0,c\geq 0$, $\sqrt{a+c}-\sqrt{b+c}\leq\sqrt{a}-\sqrt{b}$. Thus $\{[L_{i},U_{i}],i=1,\dots,N\}$ is an $u\sqrt{\frac{1-p}{1-p^{\prime}}}$-bracketing for $\mathcal{K}_{p}(A)$.
@@ B.3.2 Proof for Lemma 12
$\,\leq\rho_{0,\alpha}\left(\theta^{*},\theta\right)$
@@ B.3.3 Proof for Lemma 14
Note that $m_{n}(A_{n}^{c},\bm{X})=m_{n}(A_{n}^{c}\cap\Omega_{n},\bm{X})+m_{n}(A_{n}^{c}\cap\Omega_{n}^{c},\bm{X})$. From Lemma 9 and 10, we have
@@ B.3.4 Proof for Lemma 15
Since $f(\bm{x};\theta)$ has all directional derivatives. Let $\phi(\alpha)=\theta+\alpha(\theta^{*}-\theta)$ for some $\alpha\in(0,1)$, note that
$\log k_{p_{n}}(\bm{x};\theta)-\log k_{p_{n}}(\bm{x};\theta^{*})\leq\sup_{\alpha\in(0,1)}\|\tilde{\nabla}_{\theta}\log k_{p_{n}}(\bm{x};\phi(\alpha))\|_{2}\|\theta-\theta^{*}\|_{2}$
@@
~$\|\tilde{\nabla}_{\theta}\log k_{p_{n}}(\bm{x};\theta)\|_{2}=$ $\left\|\frac{(1-p_{n})\tilde{\nabla}_{\theta}f(\bm{x};\theta)}{(1-p_{n})f(\bm{x};\theta)+p_{n}g(\bm{x})}\right\|_{2}$
@@
Since by Assumption 4, ~$\|\tilde{\nabla}_{\theta}f(\bm{x};\theta)/g(\bm{x})\|_{2}$ is bounded as $\|\bm{w}\|_{2}\rightarrow\infty$ in all directions, we have ~$\tilde{\nabla}f(\bm{x};\theta)/g(\bm{x})$ is bounded on $\mathbb{R}^{d}\times A$ for any compact set $A\subset\Omega$. Thus there exists a sequence $T_{n}<\infty$ such that,
$\sup_{x\in\mathbb{R}^{d}}\sup_{\theta\in A_{n}}\left\|\frac{\tilde{\nabla}_{\theta}f(\bm{x};\theta)}{g(\bm{x})}\right\|_{2}\leq T_{n}$
@@ Appendix C Theoretical Application on Regression Settings
In this section, we will verify the Assumptions in Theorem 5 for linear regression, logistic regression and Cauchy regression. The analysis is presented first, followed by technical lemmas.
@@ C.1 Linear Regression Case
$\Omega_{n}:=\{(\theta,\sigma)\in\mathbb{R}^{d}:\|\theta^{*}-\theta\|_{2}\leq R_{n},\,|\sigma-\sigma^{*}|<R_{n},\sigma>1/R_{n}\}$
@@
$\left|\sqrt{f_{\theta,\sigma}(x|\bm{w})}-\sqrt{f_{\vartheta,\varsigma}(x|\bm{w})}\right|\leq M^{(n)}_{\vartheta,\varsigma,r}(x,\bm{w})\|\theta-\vartheta\|_{\infty}$
@@
where $M_{n}=\|\theta^{*}\|_{2}+R_{n}$. Then by Lemma 21, $\|M^{(n)}_{\vartheta,\varsigma,r}\|^{2}_{2}\preceq M_{n}^{12}$, $\forall(\vartheta,\varsigma)\in\Omega_{n}$ which satisfies the requirement of Lemma 3. Since $\log M_{r}^{(n)}=O(\log M_{n})$, the condition (8) is satisfied by taking
@@
for some $q>4$ and $k>0$ to be chosen for the sake of Assumptions 3.
@@
Moreover, for $R_{n}$ large, using Remark 15
@@
$M_{n}^{2}=n^{2k}\approx R_{n}^{2}\geq 4c_{1}n\phi_{n}^{2}=4c_{1}n^{(q-2)/q}$
which can be satisfied by choosing $k>(q-2)/2q$. Finally, Lemma 22 shows that the pair of sequences $\phi_{n}$ and $\psi_{n}$ in Proposition 4 can be chosen (up to a multiplicative constant) to diminish at the same rate when $(\theta,\sigma)$ is close to $(\theta^{*},\sigma^{*})$. Therefore $\psi_{n}$ can also be chosen to be $\propto\phi_{n}=n^{-1/q}$. By choosing $p_{n}$ to converge at a rate slower than $n^{-1/q}$, e.g., $p_{n}=n^{-1/2q}$, then
$\epsilon_{n}=O(n^{-1/2q}),\qquad\delta_{n}=O(\exp(-\tfrac{1}{2}c_{1}n^{1-2/q}))$
@@
where $(\theta,\sigma)\in\Omega_{n}:=\{(\theta,\sigma)\in\Omega:\|\theta-\theta^{*}\|_{2}\leq R_{n},|\sigma^{*}-\sigma|\leq R_{n},\sigma>1/R_{n}\}$, for some $R_{n}>0$. If $(\theta,\sigma),(\vartheta,\varsigma)\in\Omega_{n}$ with $\|(\theta,\sigma)-(\vartheta,\varsigma)\|_{\infty}\leq r$ then
$\left|\sqrt{f_{\theta,\sigma}(x|\bm{w})}-\sqrt{f_{\vartheta,\varsigma}(x|\bm{w})}\right|\leq M_{\vartheta,\varsigma,r}(x,\bm{w})\|(\theta,\sigma)-(\vartheta,\varsigma)\|_{\infty}$
where function $M_{\vartheta,\varsigma,r}$ is given by
$M_{\vartheta,\varsigma,r}(x,\bm{w}):=\frac{R_{n}^{2}}{2}\Bigl{(}d\|\bm{w}\|_{2}+R_{n}(|x|+M_{n}\|\bm{w}\|_{2})\Bigr{)}(|x|+M_{n}\|\bm{w}\|_{2})\\ \left(\frac{R_{n}^{2}}{2\pi}\right)^{1/4}\exp\left(-\frac{1}{4R_{n}^{2}}\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}\right)$
@@
Then, by Cauchy’s Mean Value Theorem and Cauchy-Schwarz inequality, there exists some $a\in(0,1)$, $\varphi=\vartheta+a(\theta-\vartheta)$, $\zeta=\sigma_{2}+a(\sigma_{2}-\sigma_{1})$, such that
@@
$=$ $\,\frac{x-\varphi^{\top}\bm{w}}{2\zeta^{2}}f_{\varphi,\zeta}^{\frac{1}{2}}(x|\bm{w})\,(\theta-\vartheta)^{\top}\bm{x}+\frac{1}{2}\left(\frac{(x-\varphi^{\top}\bm{w})^{2}}{\zeta^{3}}-\zeta^{-1}\right)f_{\varphi,\zeta}^{\frac{1}{2}}(x|\bm{w})\,(\sigma_{1}-\sigma_{2})$
@@
×$\,\times\left(\frac{R_{n}^{2}}{2\pi}\right)^{1/4}\exp\left(-\frac{1}{4R_{n}^{2}}\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}\right)\|(\theta,\sigma),(\vartheta,\varsigma)\|_{\infty}.$
@@
$\int_{0}^{\infty}x^{k}\frac{1}{\sqrt{2\pi R^{2}}}\exp\left(-\frac{(x-\alpha w)^{2}}{2R^{2}}\right)\mathtt{d}x\leq\beta_{k}(w)M^{k}$
@@
$=$ $\,\frac{1}{\sqrt{2\pi}}\frac{R_{n}^{5}}{4}\int_{u-}^{u_{+}}\Bigl{(}d\|\bm{w}\|_{2}+R_{n}(|x|+M_{n}\|\bm{w}\|_{2})\Bigr{)}^{2}(|x|+M_{n}\|\bm{w}\|_{2})^{2}\exp\left(-\frac{1}{2R_{n}^{2}}\right)\mathtt{d}x$
@@
Note that $\int_{-\infty}^{u_{-}}M_{\vartheta,\varsigma,r}^{2}(x,\bm{w})\mathtt{d}y$ is comprised of summing integrals of form
@@
Since $|u_{-}|,|u_{+}|\leq M_{n}\|\bm{w}\|_{2}$, the integrals $\int_{-\infty}^{u_{-}}M_{\vartheta,\varsigma,r}^{2}(x,\bm{w})\mathtt{d}y$ and $\int_{u_{-}}^{\infty}M_{\vartheta,\varsigma,r}^{2}(x,\bm{w})\mathtt{d}y$ share the same upper bound as above (see by change of variable $w=-y$). Then $\|M_{\vartheta,\varsigma,r}^{2}(x,\bm{w})\|_{2}^{2}\preceq M_{n}^{12}$.
@@
Let $\|\bm{w}\|_{\infty}\leq 1$ with non-zero entries,
@@
Then we have for $(\theta,\sigma)$ close to $(\theta^{*},\sigma^{*})$, there exists positive constant $C$ independent of $(\theta,\sigma)$ such that
$h((\theta^{*},\sigma^{*}),(\theta,\sigma))\leq\phi_{n}\implies\|(\theta^{*},\sigma^{*})-(\theta,\sigma))\|_{2}\leq C\phi_{n}$
@@
$=$ $\,f_{\varphi,\xi}^{\frac{1}{2}}(x|\bm{w})\left[\frac{x-\varphi^{\top}\bm{w}}{2\xi^{2}}(\theta-\theta^{*})^{\top}\bm{x}+\frac{1}{2}\left(\frac{(x-\varphi^{\top}\bm{w})^{2}}{\xi^{3}}-\xi^{-1}\right)(\sigma-\sigma^{*})\right].$
@@
Note that in the above integral with respect to $x$, $z$ is a standard normal random variable and $f_{\varphi,\xi}(x|\bm{w})\propto f_{z}(z)$. There exists constant $\lambda$ depending only on the (maximum) distance between $(\theta,\sigma)$ and $(\theta^{*},\sigma^{*})$ such that
@@ C.2 Logistic Regression Case
So we can apply Lemma 3 with $M^{(n)}_{\vartheta,r}(x,\bm{w})=\frac{1}{2}\|\bm{w}\|_{2}$. Hence $\|M^{(n)}_{\vartheta,r}(x,\bm{w})\|_{2}^{2}=O(d)$ is bounded above by a uniformed constant since $d$ is fixed. Thus, Assumption 1 can be satisfied with
$\phi_{n}=n^{-1/q},\quad R_{n}=n^{k}$
for some $q>4$ and $k>0$ to be chosen for Assumption 3. Again, it is not hard to check Lemma 7 holds for this likelihood and thus Assumption 2 is satisfied by applying Lemma 8 and Remark 14. Moreover, for $R_{n}$ large, by Remark 15,
@@
$M_{n}^{2}=n^{2k}\approx R_{n}^{2}\geq 4c_{1}n\phi_{n}^{2}=4c_{1}n^{(q-2)/q}$
which can be satisfied by choosing $k>(q-2)/2q$ similar to the previous example.
@@
$\epsilon_{n}=O(n^{-1/2q}),\qquad\delta_{n}=O(\exp(-\tfrac{1}{2}c_{1}n^{1-2/q}))$
@@
Let $\|\bm{w}\|_{\infty}\leq 1$ with non-zero entries,
$f_{\theta}(x|\bm{w}):=\Phi(\theta^{\top}\bm{w})^{x}\Phi(-\theta^{\top}\bm{w})^{1-x}$
where $\Phi(x)=(1+e^{-x})^{-1}$. Then we have for $\theta$ close to $\theta^{*}$, there exists positive constant $C$ independent of $\theta$ such that
$h(\theta^{*},\theta)\leq\phi_{n}\implies\|\theta^{*}-\theta\|_{2}\leq C\phi_{n}$
@@
Now, by Assumption $\|\bm{w}\|_{\infty}\leq 1$ with non-zero entries
$h^{2}(\theta^{*},\theta)=\frac{1}{8}\Phi(\varphi^{\top}\bm{w})\Phi(-\varphi^{\top}\bm{w})\left((\theta-\theta^{*})^{\top}\bm{w}\right)^{2}$

### v3

@@ B.3.1 Proof for Lemma 9
Proof.
It is enough to check that a $u$-bracket for $\mathcal{K}_{p^{\prime}}(A)$ is also a $u\sqrt{\frac{1-p}{1-p^{\prime}}}$-bracketing for $\mathcal{K}_{p}(A)$, where $1>p\geq p^{\prime}$. Suppose that $\{[L^{\prime}_{i},U^{\prime}_{i}],i=1,\dots,N\}$ is a $u$-bracketing for $\mathcal{K}_{p^{\prime}}(A)$. Pick an arbitrary $\theta\in A$, there exists $i\in\{1,\dots,N\}$ such that the $u$-bracket $L^{\prime}_{i},U^{\prime}_{i}]$ forms an envelope for $k_{p^{\prime}}(\bm{x};\theta)$, i.e.,
@@
$=$ $\,\frac{1-p}{1-p^{\prime}}u^{2},$
where the second inequality is by noting that for any $a\geq b\geq 0,c\geq 0$, $\sqrt{a+c}-\sqrt{b+c}\leq\sqrt{a}-\sqrt{b}$. Thus $\{[L_{i},U_{i}],i=1,\dots,N\}$ is a $u\sqrt{\frac{1-p}{1-p^{\prime}}}$-bracketing for $\mathcal{K}_{p}(A)$.
@@ B.3.2 Proof for Lemma 11
$\,\leq\rho_{0,\alpha}\left(\theta^{*},\theta\right).$
@@ B.3.3 Proof for Lemma 13
Note that $m_{n}(A_{n}^{c},\bm{X})=m_{n}(A_{n}^{c}\cap\Omega_{n},\bm{X})+m_{n}(A_{n}^{c}\cap\Omega_{n}^{c},\bm{X})$. From Lemmas 8 and 9, we have
@@ B.3.4 Proof for Lemma 14
Since $f(\bm{x};\theta)$ has all directional derivatives, let $\phi(\alpha)=\theta+\alpha(\theta^{*}-\theta)$ for some $\alpha\in(0,1)$. Note that
$\log k_{p_{n}}(\bm{x};\theta)-\log k_{p_{n}}(\bm{x};\theta^{*})\leq\sup_{\alpha\in(0,1)}\|\tilde{\nabla}_{\theta}\log k_{p_{n}}(\bm{x};\phi(\alpha))\|_{2}\|\theta-\theta^{*}\|_{2},$
@@
$\|\tilde{\nabla}_{\theta}\log k_{p_{n}}(\bm{x};\theta)\|_{2}=$ $\left\|\frac{(1-p_{n})\tilde{\nabla}_{\theta}f(\bm{x};\theta)}{(1-p_{n})f(\bm{x};\theta)+p_{n}g(\bm{x})}\right\|_{2}$
@@
Since by Assumption 4, $\|\tilde{\nabla}_{\theta}f(\bm{x};\theta)/g(\bm{x})\|_{2}$ is bounded as $\|\bm{w}\|_{2}\rightarrow\infty$ in all directions, hence $\tilde{\nabla}f(\bm{x};\theta)/g(\bm{x})$ is bounded on $\mathbb{R}^{d}\times A$ for any compact set $A\subset\Omega$. Thus there exists a sequence $T_{n}<\infty$ such that,
$\sup_{x\in\mathbb{R}^{d}}\sup_{\theta\in A_{n}}\left\|\frac{\tilde{\nabla}_{\theta}f(\bm{x};\theta)}{g(\bm{x})}\right\|_{2}\leq T_{n}.$
@@ Appendix C Theoretical Application on Regression Settings
In this section, we will verify the assumptions in Theorem 4 for linear regression, logistic regression, and Cauchy regression. The analysis is presented first, followed by technical lemmas.
@@ C.1 Linear Regression Case
$\Omega_{n}:=\{(\theta,\sigma)\in\mathbb{R}^{d}:\|\theta^{*}-\theta\|_{2}\leq R_{n},\,|\sigma-\sigma^{*}|<R_{n},\sigma>1/R_{n}\},$
@@
$\left|\sqrt{f_{\theta,\sigma}(x|\bm{w})}-\sqrt{f_{\vartheta,\varsigma}(x|\bm{w})}\right|\leq M^{(n)}_{\vartheta,\varsigma,r}(x,\bm{w})\|\theta-\vartheta\|_{\infty},$
@@
where $M_{n}=\|\theta^{*}\|_{2}+R_{n}$. Then by Lemma 20, $\|M^{(n)}_{\vartheta,\varsigma,r}\|^{2}_{2}\preceq M_{n}^{12}$, $\forall(\vartheta,\varsigma)\in\Omega_{n}$ which satisfies the requirements of Lemma 3. Since $\log M_{r}^{(n)}=O(\log M_{n})$, the condition (8) is satisfied by taking
@@
for some $q>4$ and $k>0$ to be chosen to satisfy Assumptions 3.
@@
Moreover, for $R_{n}$ large, by Remark 13
@@
$M_{n}^{2}=n^{2k}\approx R_{n}^{2}\geq 4c_{1}n\phi_{n}^{2}=4c_{1}n^{(q-2)/q},$
which can be satisfied by choosing k>(q2)/2qk>(q-2)/2q. Finally, Lemma 21 shows that the pair of sequences $\phi_{n}$ and $\psi_{n}$ in Proposition 3 can be chosen (up to a multiplicative constant) to diminish at the same rate when $(\theta,\sigma)$ is close to $(\theta^{*},\sigma^{*})$. Therefore $\psi_{n}$ can also be chosen to be $\propto\phi_{n}=n^{-1/q}$. By choosing $p_{n}$ to converge at a rate slower than $n^{-1/q}$, e.g., $p_{n}=n^{-1/2q}$, then
$\epsilon_{n}=O(n^{-1/2q}),\qquad\delta_{n}=O(\exp(-\tfrac{1}{2}c_{1}n^{1-2/q})),$
@@
where $(\theta,\sigma)\in\Omega_{n}:=\{(\theta,\sigma)\in\Omega:\|\theta-\theta^{*}\|_{2}\leq R_{n},|\sigma^{*}-\sigma|\leq R_{n},\sigma>1/R_{n}\}$, for some $R_{n}>0$. If $(\theta,\sigma),(\vartheta,\varsigma)\in\Omega_{n}$ with $\|(\theta,\sigma)-(\vartheta,\varsigma)\|_{\infty}\leq r$, then
$\left|\sqrt{f_{\theta,\sigma}(x|\bm{w})}-\sqrt{f_{\vartheta,\varsigma}(x|\bm{w})}\right|\leq M_{\vartheta,\varsigma,r}(x,\bm{w})\|(\theta,\sigma)-(\vartheta,\varsigma)\|_{\infty},$
where the function $M_{\vartheta,\varsigma,r}$ is given by
@@
Then, by Cauchy’s Mean Value Theorem and the Cauchy-Schwarz inequality, there exists some $a\in(0,1)$, $\varphi=\vartheta+a(\theta-\vartheta)$, $\zeta=\varsigma+a(\sigma-\varsigma)$, such that
@@
$=$ $\,\frac{x-\varphi^{\top}\bm{w}}{2\zeta^{2}}f_{\varphi,\zeta}^{\frac{1}{2}}(x|\bm{w})\,(\theta-\vartheta)^{\top}\bm{w}+\frac{1}{2}\left(\frac{(x-\varphi^{\top}\bm{w})^{2}}{\zeta^{3}}-\zeta^{-1}\right)f_{\varphi,\zeta}^{\frac{1}{2}}(x|\bm{w})\,(\sigma-\varsigma)$
@@
$\,\times\left(\frac{R_{n}^{2}}{2\pi}\right)^{1/4}\exp\left(-\frac{1}{4R_{n}^{2}}\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}\right)\|(\theta,\sigma)-(\vartheta,\varsigma)\|_{\infty}.$
@@
$\int_{0}^{\infty}x^{k}\frac{1}{\sqrt{2\pi R^{2}}}\exp\left(-\frac{(x-\alpha w)^{2}}{2R^{2}}\right)\mathtt{d}x\leq\beta_{k}(w)M^{k},$
@@
$=$ $\,\frac{1}{\sqrt{2\pi}}\frac{R_{n}^{5}}{4}\int_{u_{-}}^{u_{+}}\Bigl(d\|\bm{w}\|_{2}+R_{n}(|x|+M_{n}\|\bm{w}\|_{2})\Bigr)^{2}(|x|+M_{n}\|\bm{w}\|_{2})^{2}\exp\left(-\frac{1}{2R_{n}^{2}}\right)\mathtt{d}x$
@@
Note that $\int_{-\infty}^{u_{-}}M_{\vartheta,\varsigma,r}^{2}(x,\bm{w})\mathtt{d}x$ is comprised of summing integrals of form
@@
Since $|u_{-}|,|u_{+}|\leq M_{n}\|\bm{w}\|_{2}$, the integrals $\int_{-\infty}^{u_{-}}M_{\vartheta,\varsigma,r}^{2}(x,\bm{w})\mathtt{d}x$ and $\int_{u_{-}}^{\infty}M_{\vartheta,\varsigma,r}^{2}(x,\bm{w})\mathtt{d}x$ share the same upper bound as above (see by change of variable w=xw=-x). Then
$\|M_{\vartheta,\varsigma,r}^{2}(x,\bm{w})\|_{2}^{2}\preceq M_{n}^{12}.$
@@
Let $\|\bm{w}\|_{\infty}\leq 1$ with nonzero entries,
@@
Then we have for $(\theta,\sigma)$ close to $(\theta^{*},\sigma^{*})$, there exists a positive constant $C$ independent of $(\theta,\sigma)$ such that
$h((\theta^{*},\sigma^{*}),(\theta,\sigma))\leq\phi_{n}\implies\|(\theta^{*},\sigma^{*})-(\theta,\sigma)\|_{2}\leq C\phi_{n}.$
@@
$=$ $\,f_{\varphi,\xi}^{\frac{1}{2}}(x|\bm{w})\left[\frac{x-\varphi^{\top}\bm{w}}{2\xi^{2}}(\theta-\theta^{*})^{\top}\bm{w}+\frac{1}{2}\left(\frac{(x-\varphi^{\top}\bm{w})^{2}}{\xi^{3}}-\xi^{-1}\right)(\sigma-\sigma^{*})\right].$
@@
Note that in the above integral with respect to $x$, $z$ is a standard normal random variable and $f_{\varphi,\xi}(x|\bm{w})\propto f_{z}(z)$. There exists a constant $\lambda$ depending only on the (maximum) distance between $(\theta,\sigma)$ and $(\theta^{*},\sigma^{*})$ such that
@@ C.2 Logistic Regression Case
So we can apply Lemma 2 with $M^{(n)}_{\vartheta,r}(x,\bm{w})=\frac{1}{2}\|\bm{w}\|_{2}$. Hence $\|M^{(n)}_{\vartheta,r}(x,\bm{w})\|_{2}^{2}=O(d)$ is bounded above by a uniform constant since $d$ is fixed. Thus, Assumption 1 can be satisfied with
$\phi_{n}=n^{-1/q},\quad R_{n}=n^{k},$
for some $q>4$ and $k>0$ to be chosen to satisfy Assumption 3. Again, it is not hard to check Lemma 6 holds for this likelihood and thus Assumption 2 is satisfied by applying Lemma 7 and Remark 12. Moreover, for $R_{n}$ large, by Remark 13,
@@
$M_{n}^{2}=n^{2k}\approx R_{n}^{2}\geq 4c_{1}n\phi_{n}^{2}=4c_{1}n^{(q-2)/q},$
which can be satisfied by choosing k>(q2)/2qk>(q-2)/2q, similar to the previous example.
@@
$\epsilon_{n}=O(n^{-1/2q}),\qquad\delta_{n}=O(\exp(-\tfrac{1}{2}c_{1}n^{1-2/q})),$
@@
Let $\|\bm{w}\|_{\infty}\leq 1$ with nonzero entries,
$f_{\theta}(x|\bm{w}):=\Phi(\theta^{\top}\bm{w})^{x}\Phi(-\theta^{\top}\bm{w})^{1-x},$
where $\Phi(x)=(1+e^{-x})^{-1}$. Then we have for $\theta$ close to $\theta^{*}$, there exists a positive constant $C$ independent of $\theta$ such that
@@
Now, by Assumption $\|\bm{w}\|_{\infty}\leq 1$ with nonzero entries
$h^{2}(\theta^{*},\theta)=\frac{1}{8}\Phi(\varphi^{\top}\bm{w})\Phi(-\varphi^{\top}\bm{w})\left((\theta-\theta^{*})^{\top}\bm{w}\right)^{2}.$